# Principal-Agent Multitasking: the Uniformity of Optimal Contracts and its Efficient Learning via Instrumental Regression

### v1

@@ Abstract
This work studies the multitasking principal-agent problem. I first show a “uniformity” result. Specifically, when the tasks are perfect substitutes, and the agent’s cost function is homogeneous to a certain degree, then the optimal contract only depends on the marginal utility of each task and the degree of homogeneity. I then study a setting where the marginal utility of each task is unknown so that the optimal contract must be learned or estimated with observational data. I identify this problem as a regression problem with measurement errors and observe that this problem can be cast as an instrumental regression problem. The current works observe that both the contract and the repeated observations (when available) can act as valid instrumental variables, and propose using the generalized method of moments estimator to compute an approximately optimal contract from offline data. I also study an online setting and show how the optimal contract can be efficiently learned in an online fashion using the two estimators. Here the principal faces an exploration-exploitation tradeoff: she must experiment with new contracts and observe their outcome whilst at the same time ensuring her experimentations are not deviating too much from the optimal contract. This work shows when repeated observations are available and agents are sufficiently “diverse", the principal can achieve a very low $\widetilde{O}(d)$ cumulative utility loss, even with a “pure exploitation" algorithm.
@@ 1 Introduction
A firm’s owner (the principal) hires a manager (the agent) to manage the firm. The manager is responsible for several tasks, for example, managing personnel, maintaining a healthy work environment, planning and executing activities, etc. The manager can allocate a certain effort to each task, however, this effort is unobservable and unverifiable to the owner of the firm. Instead, the owner can observe a noisy signal, indicating the performance of each task. The principal can also observe the total revenue generated by the firm under the manager’s management, which will depend on the effort the manager puts into the tasks. The owner of the firm must incentivize the manager to exert effort through contracts. Since the exact effort is unobservable, the contract can only depend on the noisy signals associated with each task. Such is the basic model for the multitask principal-agent problem, which was first studied in the seminal work [1].
@@
It often happens that the principal contracts with more than one agent. For example, in franchising scenarios, the principal is the franchiser and the agents are franchisees ([2]). The franchiser must design contracts to incentivize each franchisee to exert the appropriate amount of effort to generate revenue. Another example is e-commerce platforms, such as YouTube. YouTube is a video-sharing platform where content creators earn revenue through advertisements. According to this documentation ([3]), YouTube content creators earn 55% of the total advertisement revenue, while the remaining 45% goes to YouTube. While intuition might suggest that the principal should tailor her contracts when facing different agents to maximize her utility, in the above two scenarios, it is often observed that the contract is uniform across different agents. This leads to our first research question.
@@
How can the uniformity of the contract terms be justified in (multitask) principal-agent problems?
@@
As an answer to this question, I show that when the agent’s cost function is homogeneous to a certain degree and the principal’s utility takes a linear form across tasks (i.e., tasks are perfect substitutes), then the optimal contract is indeed uniform for different agents. In particular, the optimal contract only depends on the marginal utility of each task and the degree of homogeneity of the agent’s cost function. The implications of this result are practically very attractive. It implies that there is no need for the principal to discriminate against agents in an attempt to maximize her own utility, i.e., the optimal contract is fair. The result here follows closely in spirit to a result in the important work by [2]; the difference is that in their work they study a single task setting with double moral hazard while this work studies a multitask setting.
@@
The above result shows that in the optimal contract, the principal should transfer more utility to the agent for more “important” tasks, which aligns with intuition. However, it may be the case that the exact marginal utility of each task is unknown. In this case, the principal must learn this value with observational data in order to design the optimal contract. This leads to our second research question.
@@
Assuming again that the principal’s utility takes a linear form across tasks, estimating the marginal utilities of each task is essentially a linear regression problem, where the total revenue is the response variable, and the observed signals of each task are the covariates. However, since the observed signals are noisy measurements of the true effort, the regression problem has a measurement error problem ([4]). This work identifies the contract as a valid instrumental variable and proposes estimation and learning algorithms based on the generalized method of moments (GMM) estimator ([5]). Moreover, when repeated observations are available, the GMM estimator based on repeated observations achieves even faster convergence when the agents are “diverse”.
It seems important to point out that a problem setting very similar to the current work was studied in the (fairly) recent work [6]. Their work studied optimization algorithms in Stackelberg games from revealed preferences. As a special case, they gave a discussion on how their algorithm can be applied to the multitask principal-agent problem (though their work did not explicitly identify their problem as such). The problem of learning the optimal contract is also studied in this work, however, the approach is fundamentally different from theirs. This is because the current work also identified a “uniformity” result, which seemed to escape their analysis. The problem of estimating and learning the optimal contract in this work uses instrumental regression with the generalized method of moments method and is based on this “uniformity” result, whereas in their work the problem is solved via a two-stage optimization algorithm.
2 Related Work
@@
Contract theory studies a principal-agent problem where the principal must incentivize the agent to exert effort through contracts. Some early important works in the economics community include [7, 8, 1]. The work by [1] studies the multitask principal-agent problem, which serves as the starting point of the current work. Recently, computational aspects of contract design have been studied by many works in the computer science community. For example, [9, 10, 11] study combinatorial algorithms in contract design; [12, 13, 14] study learning algorithms for contract design.
@@
Measurement error model is a well-studied problem in econometrics and statistics, for a textbook treatment the reader is referred to [4]. The instrumental regression and generalized method of moments (GMM) estimators (again for a textbook treatment see [5]) are popular methods to address endogeneity issues in models, including the measurement error model. Some recent works in computer science also notice endogeneity issues in learning/classification problems and apply the instrumental regression method in the context of strategic classification ([15]), online learning/bandits ([16]), reinforcement learning ([17]), etc.
@@
Bandit problems are a type of online learning problem with partial feedback. Typically, bandit algorithms need to balance exploration (experimenting with unknown actions) and exploitation (choosing actions whose reward is estimated to be higher). Research on bandit problems is too broad to cover here, but for an overview see [18]. Recently, a number of papers have sought to understand the performance of the pure exploitation (i.e., greedy) algorithm in bandit problems. In particular, for linear bandits, the success of the greedy algorithm is explained under the framework of smoothed analysis [19, 20, 21].
3 Multitask Principal-Agent: Problem Description and the Uniformity Result
I study a multitasking principal-agent problem. The agent can exert effort in $d$ dimensions, where each dimension can be interpreted as a task. Denote effort of the agent by the vector $w\in(\mathbb{R}^{d})^{+}$, and assume that the agent incurs a private cost $c(w)\in\mathbb{R}^{+}$ when exerting effort $w$. The exact effort is unobservable and unverifiable to the principal, but a signal $x\in\mathbb{R}^{d}$ is generated and observed by the principal. The signal is assumed to satisfy
$\mathbb{E}[x|w]=w.$
@@ 3 Multitask Principal-Agent: Problem Description and the Uniformity Result
The agent’s effort gives a private benefit $y$ to the principal, and it is assumed this benefit depends on the agents true efforts through the linear model:
$\mathbb{E}[y|w]=\langle\theta,w\rangle.$
Here $\theta\in(\mathbb{R}^{d})^{+}$. The value $\theta_{i}$ can be interpreted as the marginal utility of task $i$.
The principal incentivizes the agent to exert effort through a contract $\beta\in\mathbb{R}^{d}$. Under this contract, the monetary transfer to the agent is $\langle\beta,x\rangle$ when the realized signal is $x$. The agent then takes the action $w=w(\beta)$ which maximizes his expected utility under the contract $\beta$:
$w=\arg\max_{w^{\prime}}\langle\beta,w^{\prime}\rangle-c(w^{\prime}).$
Then, under the contract $\beta$, assuming the agent takes his best action $w$, the principal’s expected utility is
$\langle\theta,w\rangle-\langle\beta,w\rangle,$
i.e., the expected private benefit minus the expected payment to the agent. The optimal contract is the contract maximizing the principal’s expected utility.

### v2

# Linear Contracts in Multitasking: Robustness, Uniformity, and Learning

@@ Abstract
In this work, we study the multitasking principal-agent problem. The agent performs several task for the principal, and the principal posts a contract incentivizing the agent to exert effort. The principal can observe a signal for each task, and the contract is a mapping from the space of possible signals to a payment. We study the special class of linear contracts from three perspectives: robustness, uniformity, and learning. Firstly, we show a robustness result: in an ambiguous setting when only first moment information is known, there is a linear contract maximizing the principal’s payoff in a worst-case scenario. Secondly, we show a uniformity result: when the agent’s cost function is homogeneous to a certain degree and the the principal’s utility takes a linear form across tasks, then the optimal contract depends on the agent’s cost function only through its homogeneuity degree. Thirdly, we study the problem of learning an optimal linear contract through observational data. We identify this as an measurement error model, and propose instrumental regression methods to estimate the optimal contract parameters in an offline setting, or to learn the optimal contract in an online setting.
@@ 1 Introduction
A firm’s owner (the principal) hires a manager (the agent) to perform multiple tasks, such as managing personnel, maintaining a healthy work environment, and executing strategic initiatives. The agent can allocate effort across these tasks, but this effort is unobservable and unverifiable to the principal. Instead, the principal receives noisy signals that indicate task performance and can observe the total revenue generated by the firm, which depends on the agent’s effort allocation. To incentivize the agent, the principal offers a contract that links compensation to these noisy signals. Since direct effort observation is impossible, the contract structure plays a crucial role in aligning incentives. This setting defines the multitask principal-agent problem, first studied in the seminal work [Holmstrom and Milgrom, 1991].
A contract in a multitasking problem is a mapping from the signal space to a payment. In theory, such contracts can be highly complex. Even in the well-studied single-task setting, optimal contracts can exhibit significant complexity. However, empirical observations suggest that most contracts used in practice tend to be relatively simple. A particularly important class of simple contracts is linear contracts. Many recent works have examined linear contracts in single task settings from various perspectives, including robustness and approximation properties (e.g. [Dütting et al., 2019], [Carroll, 2015]). However, the role of linear contracts in multitask settings remains underexplored. In this work, we study linear contracts in multitask principal-agent problems from three key perspectives: robustness, uniformity, and learnability.
Robustness.
Recent works have studied the robustness of linear contracts and established their optimality in ambiguous or non-Bayesian worst-case settings (e.g. [Carroll, 2015], [Dütting et al., 2019]). In this work, we extend these findings to the multitask principal-agent setting and provide a robustness guarantee under ambiguity. Our research question can be summarized as follows:
Does the linear contract exhibit strong robustness guarantees in the multitask principal-agent setting?
We show that when the principal knows only the first moment of each agent’s action but lacks information about its full distribution, a linear contract remains optimal under the worst-case distribution. This result highlights the structural advantages of linear contracts in environments with uncertainty. Our findings are similar in spirit to the prior results in [Dütting et al., 2019], which demonstrate the robustness of linear contracts in single-task settings. However, the transition to multitask environments introduces additional complexities, requiring new proof ideas; the crux of our proof contains a triangulation result from convex geometry followed by an application of Kakutani’s fixed-point theorem.
Uniformity.
@@
In many real-world scenarios, the principal engages with multiple agents under the same contractual framework. A prominent example is franchising, where the principal (franchiser) contracts with multiple franchisees to ensure consistent effort and revenue generation [Bhattacharyya and Lafontaine, 1995]. Another example is e-commerce platforms such as YouTube, where content creators earn a share of advertising revenue. According to official documentation [Google, 2024], YouTube content creators receive 55% of the total advertisement revenue, while the remaining 45% is retained by the platform. While one might expect the principal to tailor contracts for different agents to maximize utility, it is often observed that contracts remain uniform across agents. This observation leads to our second research question.
@@
How can the uniformity of contract terms be justified in multitask principal-agent problems?
@@
To answer this question, we show that when the agent’s cost function exhibits a certain degree of homogeneity and the principal’s utility is linear across tasks, the optimal contract is necessarily uniform across agents. Specifically, the optimal contract depends on the agent’s cost function only through it’s the homogeneity degree. This result has strong practical implications: it suggests that the principal does not need to differentiate contracts to maximize utility, reinforcing fairness in contract design. Our findings are similar in spirit to the work of [Bhattacharyya and Lafontaine, 1995], who examine a single-task setting with double moral hazard. In contrast, our study focuses on multitask environments.
Learning
@@
In the optimal contract, imaginably the principal should offer a higher incentives to tasks that are more “important”, i.e., tasks which have a higher marginal utility. However, the exact “importance” of each task may not always be known. In such cases, the principal must infer these values from observational data to design the optimal contract. This naturally leads to our third research question:
@@
Assuming that the principal’s utility remains linear across tasks, estimating the marginal utilities reduces to a linear regression problem where her realized private benefit serves as the response variable and the observed signals for each task act as covariates. However, the observed signals are noisy representations of the agent’s true effort, leading to a measurement error model [Fuller, 2009]. In classical regression, measurement error in covariates introduces bias, making naive least squares estimators inconsistent and unreliable. This bias arises because the noise in observed signals weakens the correlation between true effort and output, leading to attenuation effects where coefficient estimates are systematically biased toward zero.
To address this issue, we leverage instrumental variable techniques, identifying the contract itself as a valid instrument. Specifically, we propose estimation and learning algorithms based on the generalized method of moments (GMM) estimator [Greene, 2008]. The GMM approach corrects for the endogeneity introduced by measurement error by exploiting moment conditions that remain valid despite the noise. Furthermore, when repeated observations are available, they can serve as additional instrumental variables, enhancing estimation accuracy. The effectiveness of this method is particularly pronounced when agents exhibit sufficient diversity in their responses, facilitating a faster convergence rate to the true parameter. Specifically, we show that the principal can achieve a $\widetilde{O}(d)$ regret in utility loss, even adopting a pure exploitation-type algorithm.
2 Additional Related Work
@@
Contract theory studies a principal-agent problem where the principal must incentivize the agent to exert effort through contracts. Some early important works in the economics community include [Holmström, 1979, Holmstrom, 1982, Holmstrom and Milgrom, 1991]. The work by [Holmstrom and Milgrom, 1991] studies the multitask principal-agent problem, which serves as the starting point of the current work. Recently, computational aspects of contract design have been studied by many works in the computer science community. For example, [Dütting et al., 2019, Duetting et al., 2024, Dütting et al., 2022] study combinatorial algorithms in contract design; [Ho et al., 2014, Zuo, 2024, Guruganesh et al., 2024] study learning algorithms for contract design.
@@
Measurement error model is a well-studied problem in econometrics and statistics, for a textbook treatment the reader is referred to [Fuller, 2009]. The instrumental regression and generalized method of moments (GMM) estimators (again for a textbook treatment see [Greene, 2008]) are popular methods to address endogeneity issues in models, including the measurement error model. Some recent works in computer science also notice endogeneity issues in learning/classification problems and apply the instrumental regression method in the context of strategic classification ([Harris et al., 2022]), online learning/bandits ([Della Vecchia and Basu, 2023]), reinforcement learning ([Yu et al., 2022]), etc.
@@
Bandit problems are a type of online learning problem with partial feedback. Typically, bandit algorithms need to balance exploration (experimenting with unknown actions) and exploitation (choosing actions whose reward is estimated to be higher). Research on bandit problems is too broad to cover here, but for an overview see [Lattimore and Szepesvári, 2020]. Recently, a number of papers have sought to understand the performance of the pure exploitation (i.e., greedy) algorithm in bandit problems. In particular, for linear bandits, the success of the greedy algorithm is explained under the framework of smoothed analysis [Bastani et al., 2021, Kannan et al., 2018, Sivakumar et al., 2022].
3 The Model
We study the multitasking principal-agent problem, where an agent exerts effort across $d$ dimensions, each representing a distinct task. Let the agent’s effort be denoted by the vector $a\in\mathcal{A}$, where $\mathcal{A}$ represents the set of feasible actions available to the agent. Depending on the setting, $\mathcal{A}$ may be $(\mathbb{R}^{+})^{d}$ or a subset thereof. For clarity, we specify assumptions on $\mathcal{A}$ as needed in each section.
The agent incurs a private cost $c(a)\in\mathbb{R}^{+}$ when exerting the effort vector $a$. We will always assume $c(a)$ is strictly increasing, strictly convex, and twice differentiable. While the principal cannot directly observe or verify the exact effort level $a$, she receives a signal $x\in\mathbb{R}^{d}$, which serves as an imperfect but unbiased indicator of the agent’s effort:
$\mathbb{E}[x|a]=a.$
@@ 3 The Model
The principal incentivizes the agent to exert effort through a contract. We focus primarily on linear contracts, parameterized by $\beta\in(\mathbb{R}^{+})^{d}$. Under such a contract, the agent receives a monetary transfer of $\langle\beta,x\rangle$ upon the realization of signal $x$. The agent, in turn, selects an action $a=a(\beta)$ that maximizes their expected utility given the contract terms:
$a=\arg\max_{a^{\prime}}\langle\beta,a^{\prime}\rangle-c(a^{\prime}).$
The agent’s effort $a$ gives a noisy private benefit $y(a)$ to the principal, and the principal’s expected payoff, or expected utility, is the expected private benefit minus the expected payment to the agent. In the expression below, letting $a$ be the best response to the contract $\beta$, the principal’s expected utility is then
$u(\beta)=\mathbb{E}[y(a)]-\langle\beta,a\rangle.$
To sum up, the interaction protocol can be summarized as follows.
  1. 1.
    The principal posts a contract $\beta\in\mathbb{R}^{d}$
  2. 2.
    The agent best responds with a private effort (action) $a\in\mathbb{R}$, and suffers private cost $c(a)$
  3. 3.
    A noisy signal $x$ is generated and that $\mathbb{E}[x|a]=a$
  4. 4.
    The agent’s expected utility is $\langle\beta,a\rangle-c(a)$
  5. 5.
    The principal’s expected utility is $\mathbb{E}[y(a)]-\langle\beta,a\rangle$.
Affine Contracts
In certain parts of the proof we will also consider affine contracts, which are linear contracts with the addition of an offset term. Under the affine contract $\gamma\in\mathbb{R}^{d+1}$, the payment to the agent (abusing notation) is $\gamma(x)=\gamma_{0}+\sum_{i\in[d]}\gamma_{i}x_{i}$.
A Preliminary Lemma
Before going into our technical results, we first give a preliminary lemma.
Lemma 1 (Lemma 9 in [Roth et al., 2016]).
Suppose the agent’s action set is downward-closed, meaning if $a\in\mathcal{A}$, then $\rho a\in\mathcal{A}$ for some $\rho\in(0,1)$. Among all non-negative linear or affine contracts that induce some non-zero action $a$, the linear contract $\beta=\nabla c(a)$ induces $a$ at least expected payment.
4 Robustness in Ambiguous settings
Here, we explore the robustness properties of linear contracts in an ambiguous setting. In particular, we show that linear contracts are worst-case optimal in environments where only the mean outcomes of each action are known.
[…]
The principal seeks to guarantee a worst-case payoff in an ambiguous setting where the exact distribution of each action is unknown. To formally define the worst-case payoff, we first introduce the concept of compatible distributions.
@@ 4 Robustness in Ambiguous settings
[…]
$\mathbb{E}_{x\sim F_{a}}[x].$
A contract by a principal is a mapping from all possible outcomes to a payment, which we denote by $w:\mathcal{X}\rightarrow\mathbb{R}^{+}$. Note that the agent is protect by limited liability. The worst-case payoff for the principal in an ambiguous setting is defined as
$u(w)=\min_{\mathcal{F}\text{ is compatible}}u(w;\mathcal{F})$ (1)
Here $u(w;\mathcal{F})$ is the principal’s payoff using contract $w$ assuming the agent’s true distributions were $\mathcal{F}$ and that the agent best responds to the contract $w$.
We show that there is a linear contract $\beta\in\mathbb{R}^{d}$ maximizing the above expression. Recall that under the linear contract $\beta$, the agent’s payment under signal $x$ is
$\beta(x)=\sum_{i=1}^{d}\beta_{i}x_{i}.$
Theorem 1.
There exists a linear contract maximizing the principal’s worst-case objective as given in eq. 1.
Remark 1.
[…]
4.1 Proof Sketch
Let’s first focus on affine contracts. We show that there exists an affine contract that maximizes the principal’s worst-case objective. A key property of affine contracts is that under any family of compatible distributions over outcomes, the agent’s best response remains unchanged—so the principal’s payoff is independent of the particular distribution in that family. Then in the last step, we improve the affine contract to a linear contract.
Our strategy consists of the following steps.
From a Non-affine Contract to an Affine Contract.
Given any (possibly non-affine) contract $w(\cdot)$, we will construct an affine contract $\phi$ and identify a corresponding family of compatible distributions $\mathcal{F}$ such that the principal’s payoff under $\phi$ is at least as high as under $w$ for the compatible distribution family $\mathcal{F}$. Since the payoff from an affine contract does not depend on the particular choice of distribution family $\mathcal{F}$, it follows that the worst-case objective achieved by some affine contract is at least as good as that achieved by $w(\cdot)$.
A Family of Affine Contract via Upper Supporting Hyperplanes.
To obtain the affine contract, we upper-bound $w(\cdot)$ by finding an affine function $\phi$ that satisfies
$\phi(x)\geq w(x)\quad\text{for all }x.$
Consider the lifted set
$S=\{(x,w(x)):x\in\mathcal{X}\}.$
For any hyperplane $H$ that supports $S$ from above (an upper supporting hyperplane), we can associate an affine contract $\phi(H)$ that satisfies the inequality above. In this way, $\phi(H)$ serves as an upper bound to the payment function $w$.
“Self-Owned” Hyperplanes.
We now introduce the notion of a self-owned hyperplane: a hyperplane $H$ is called self-owned if the induced action $a(H)$ under the affine contract $\phi(H)$ lies in the convex hull of its own contact set (after projected back to $\mathcal{X}$), that is,
$a(H)\in\operatorname{conv}(P(H)).$
Here we define
$P(H)=\{x\in\mathcal{X}:(x,w(x))\in H\}.$
When $H$ is self-owned, one may choose a compatible distribution where the support of taking action $a(H)$ is only supported on $P(H)$. Then, since $H$ touches the $S$, we know that the agent’s payment is identical under both $w$ and $\phi(H)$. This means that if the affine contract $\phi(H)$ induced action $a$, then under this compatible distribution the non-affine contract $w(\cdot)$ also induces action $a$, and that $\phi(H)$ induces $a$ at a no larger payment.
Existence of a Self-Owned Hyperplane.
A classical result in convex geometry (e.g. [Lee and Santos, 2017]) implies that the union of the convex hulls of $P(H)$ over all upper supporting hyperplanes $H$ covers the entire outcome space $\mathcal{X}$. In particular, for every action $a\in\mathcal{X}$ there exists some hyperplane $H$ such that $a\in\operatorname{conv}(P(H))$. Then, we can construct a set-valued mapping:
$g:\mathcal{H}\to 2^{\mathcal{H}},$
defined as
[…]
By applying Kakutani’s fixed point theorem ([Kakutani, 1941]), there exists a hyperplane $H^{*}$ such that
$H^{*}\in g(H^{*}).$
Then we have precisely found a self-owned hyperplane.
Refining to a Linear Contract.
Finally, we show that if the affine contract $\phi$ induces an action $a^{*}$, then there exists a linear contract (i.e., one with zero offset) that also induces $a^{*}$ while providing no greater payment to the agent. This completes the proof by establishing that within the class of contracts, a linear contract exists that maximizes the principal’s worst-case objective.
By these steps, we conclude that for any arbitrary contract $w$, there is a linear contract that performs at least as well in the principal’s worst-case evaluation. This completes the proof sketch.
4.2 An Example
We illustrate the key ideas, in particular, the existence of a self-owned hyperplane, with a concrete example.
Consider a setting with $d=2$ and the following non-affine (non-linear) contract $w$ defined on $\mathcal{X}=\{0,1\}^{2}$:
$w(0,0)=0,\quad w(1,0)=2,\quad w(0,1)=2,\quad w(1,1)=3.$
We lift the space $\mathcal{X}$ into $\mathbb{R}^{3}$ by appending the contract value as a third coordinate. In other words, we consider the set
$S=\{(0,0,0),\ (1,0,2),\ (0,1,2),\ (1,1,3)\}.$
If $w$ were affine, then $S$ would lie on a single hyperplane in $\mathbb{R}^{3}$. Since $w$ is not affine, $S$ forms a polytope with several facets. In particular, there are two upper facets of $S$ that yield upper supporting hyperplanes, each corresponding to an affine contract that upper bounds $w$.
One such upper facet defines an affine contract $\phi_{1}$ with
$\phi_{1}(0,0)=0,\quad\phi_{1}(1,0)=2,\quad\phi_{1}(0,1)=2,\quad\phi_{1}(1,1)=4.$
This hyperplane touches $S$ at the points
$X_{1}=\{(0,0),\ (1,0),\ (0,1)\}.$
Similarly, another upper facet defines an affine contract $\phi_{2}$ given by
$\phi_{2}(0,0)=1,\quad\phi_{2}(1,0)=2,\quad\phi_{2}(0,1)=2,\quad\phi_{2}(1,1)=3,$
supported on the points
$X_{2}=\{(1,0),\ (0,1),\ (1,1)\}.$
Note the triangulation is clearly illustrated: $\mathcal{X}=\mathtt{conv}(X_{1})\cup\mathtt{conv}(X_{2})$.
Besides the two upper facets, any convex interpolation between these two hyperplanes also yields a valid supporting hyperplane for $S$. Such interpolated hyperplanes are supported on the two endpoints $(1,0)$ and $(0,1)$.
We now claim that there must exist at least one hyperplane among these—one that we call self-owned—for which the induced action lies within its own set of support points (after projecting back to $\mathcal{X}$). In other words, if $H$ is the hyperplane corresponding to an affine contract $\phi(H)$, define
$P(H)=\{x\in\mathcal{X}\mid(x,w(x))\in H\}.$
Then $H$ is self-owned if the action $a$ induced by $\phi(H)$ satisfies
$a\in\operatorname{conv}(P(H)).$
Suppose, for the sake of contradiction, that no self-owned hyperplane exists. Then the affine contract $\phi_{1}$ would induce an action falling in $X_{2}$ while $\phi_{2}$ would induce an action in $X_{1}$. However, the mapping from the hyperplane to the induced action is continuous (small changes in the hyperplane lead to small changes in the induced action). Therefore, as one continuously interpolates between $\phi_{1}$ and $\phi_{2}$, the induced action must eventually lie exactly in the convex combination (or “interval”) determined by $(1,0)$ and $(0,1)$. Note that here Kakutani’s fixed point theorem reduces to an intermediate value theorem. Hence, this guarantees the existence of a self-owned hyperplane.
5 Uniformity of the Optimal Linear Contract
In this section, we give a characterization of the optimal linear contract. Specifically, we present a uniformity result, which states that if the agent’s cost function exhibits a certain degree of homogeneity, then the optimal linear contract depends on the cost function only through its homogeneity degree. From this point onward, we assume that the agent’s action space is $\mathcal{A}=(\mathbb{R}^{+})^{d}$.
We first introduce several standard definitions and outline the conditions imposed on the problem instance. Throughout the following, the following assumption will be imposed on the cost function of the agent(s).
Assumption 1.